# Galactic transient sources with the Cherenkov Telescope Array

### v1

@@ 4.1.1 SS 433
Abeysekara et al. (2018a) reported the results of the High Altitude Water Cherenkov (HAWC) Observatory of SS 433, collected between November 2014 and December 2017. In 1017 days of measurements with HAWC, an excess of gamma rays with a post-trial significance of 5.4 has been observed in a joint fit of the eastern and western interaction regions (namely e1 and w1) of the jets of SS 433. The quality of the data allow us to set only ULs on the angular size of the emission regions, namely 0.25∘ for the east hotspot and 0.35∘ for the west hotspot at 90% confidence level. The VHE gamma-ray flux is consistent with a hard E-2 spectrum, though current TeV flux data from HAWC are not of sufficient significance to constrain the spectral index. The reported HAWC flux at 20 TeV is 2.4 TeV-1 cm-2 s-1 at e1, and 2.1 TeV-1 cm-2 s-1 at w1. HAWC detects VHE emission from the interaction regions up to at least 25 TeV, which is the highest energy measured up to now from microquasars.
The SS 433 region was also observed by H.E.S.S. as part of a Galactic plane survey, and during two further dedicated campaigns, resulting in about 200h of observations (H. E. S. S. Collaboration et al., 2024). These observations confirmed presence of two regions of gamma-ray emission at the positions of the eastern and western jets, with peak statistical significances of 7.8 and 6.8, respectively. The reported fluxes at 1 TeV (( TeV-1 cm-2 s-1 and ( TeV-1 cm-2 s-1 for the eastern and western jets correspondingly) are inline with HAWC data taking into account the spectral slopes of and . Quality of the H.E.S.S. data also allow to study the energy dependence of the source morphology, demonstrating that while the gamma-ray emission above 10 TeV appears only at the base of the jets, the lower-energy gamma rays have their peak surface brightnesses at locations further along each jet, reflecting an energy-dependent particle energy loss timescale.
@@
Kimura et al. (2020) investigated both hadronic and leptonic scenarios for the GeV-TeV gamma-ray emission from SS 433. They found that the existing data can be explained by both the hadronic and leptonic models. The hadronic scenarios predict higher-energy photons than the leptonic scenarios, and future observations with CTA and LHAASO are needed to unravel the emission mechanism of GeV-TeV gamma rays. A preliminary association of LHAASO sources with known TeV sources based on their position reported in the first LHAASO catalog of gamma-ray sources (Cao et al., 2023b) indicate the presence of 2 LHAASO sources located at 0.26∘ and 0.15∘ from the western and eastern lobes of SS 433 correspondingly, but further observations are needed to confirm these associations.
To test the capability of CTA to study the origin of the VHE emission from the source we simulate the source with both the east and west lobes as extended sources in the energy range 0.1 - 100.0 TeV. We simulate observations with 10, 30 and 50 hours of exposure time. These simulations were repeated on each of the lobes assuming different extensions roughly corresponding to the different characteristic scales of source morphology as observed with H.E.S.S. (H. E. S. S. Collaboration et al., 2024) (see Table 3), to provide the expected detectability in terms of their size. The TS and flux of each lobe was measured for both the Northern and Southern CTA arrays, see Tables 4 and 5.
Model | ||
1 | 0.001 | 0.001 |
2 | 0.0625 | 0.0875 |
3 | 0.175 | 0.125 |
4 | 0.26 | 0.1875 |
Model | ||||||
1 | 52.0 | 76.8 | 229.1 | 303.7 | 354.3 | 477.2 |
2 | 38.9 | 10.9 | 121.9 | 105.2 | 152.0 | 114.5 |
3 | 7.9 | 9.4 | 23.7 | 57.6 | 57.1 | 100.2 |
4 | 2.4 | 8.2 | 6.3 | 20.1 | 9.4 | 37.5 |
Model | ||||||
1 | 137.4 | 184.6 | 738.8 | 821.9 | 886.3 | 1261.4 |
2 | 59.2 | 73.6 | 178.4 | 384.2 | 273.1 | 401.0 |
3 | 11.1 | 31.3 | 76.3 | 104.3 | 79.5 | 226.7 |
4 | 2.6 | 7.2 | 44.4 | 60.3 | 34.7 | 115.9 |
Central source emission and putative variability
While we do not expect to detect variability in the VHE emission coming from the lobes, microquasars are known to have flaring emission on various timescales coming from its central source. To test the possibility of CTA to detect a central source and its putative variability we simulate the local region of SS 433 with the diffuse background and the nearby MGRO 1906+06 source, where SS 433 consists of both the aforementioned lobes and a central point source. All simulations were performed when the H.E.S.S. ULs on the central source flux were not published yet and we assumed that the flux of this central source is approximately half that of the western lobe, which is roughly consistent with the ULs reported in (H. E. S. S. Collaboration et al., 2024). The lobes and the central source were both assumed to have an index of -2.0 and a pivot energy of 20 TeV (see Abeysekara et al., 2018a). The lobes were given an extension that matches Model 2 as described in Table 3. Batches of 100 simulations were run using the North and South site IRFs at multiple different exposure times ranging from 30 minutes to 50 hours. The flux was calculated after every simulation and the total data was compiled into histograms for each exposure time to determine the error range of the detections. As SS 433 could be viewed from both the north and south hemispheres the results of error measurements from both sites were compared to determine which can produce more sensitive detections. Fig. 5 shows the the comparison of this flux error ratio. Obtained results demonstrate that Southern array is better suited to study the source.
As previously mentioned, the region surrounding SS 433 does not contain many other TeV sources and could be described as an uncrowded region. Conversely, sources such as the gamma-ray binary LS 5039 are located in a region with multiple nearby TeV sources (crowded region). Using the previous setup, simulations were run for a source modelled like the central SS 433 source located at the position of LS 5039 to determine if a crowded region would change the results. It was found that the observed flux of the central source and its error were very similar for both cases of a crowded and uncrowded regions.
@@
The procedure used to test the effectiveness of CTA on dim transient sources can also be tested with sources of arbitrary intensity to determine what level of variability can be observed with shorter exposures. The results of observing sources in the same position and region of SS 433 with different fluxes ( TeV) are shown in Fig. 6. Similarly to the simulations of the SS 433 central source at different exposures (Fig 5), the Southern array configuration consistently produce a lower ratio between the 1 sigma error and the input flux of the simulation. Based on the simulations run at different fluxes, sources with a photon flux ph cm-2 s-1 will require exposure times of more than 10 hours in order to detect any expected variability. However, for sources with a flux ph cm-2 s-1, CTA may be able to detect variability as low as observing from 5 to 10 hours. At a photon flux ph cm-2 s-1 the 1 sigma ratio gets as low as with exposure times of hour long, meaning that even low variability may be detectable from relatively bright sources with very short observations.
@@ 4.1.2 Cyg X-3
Cyg X-3 is a HMXB located at a distance of 9 kpc (Reid & Miller-Jones, 2023). The companion star is a Wolf-Rayet (WR) with a strong wind mainly composed of helium. The nature of the compact object is still unknown, although a black hole scenario is favored (Zdziarski et al., 2013; Antokhin et al., 2022). The orbital period is very short, hours, indicating that the compact object is very close to the WR star, totally enshrouded in its stellar wind (orbital distance cm). The binary system is known to produce giant radio flares (flux > 10 Jy), produced by synchrotron processes from a relativistic jet oriented very close to the line of sight. Transient gamma-ray activity above 100 MeV was reported for the first time in 2009 by AGILE (Tavani et al., 2009) and Fermi-LAT (Fermi LAT Collaboration et al., 2009). The flaring activity (typical duration: 1-2 days) was observed in coincidence with a repetitive pattern of multi-frequency emission (Piano et al., 2012): the gamma-ray flares have been detected (i) during soft X-ray spectral states (around minima of the hard X-ray light curve), (ii) in the proximity of spectral transitions, and (iii) a few days before giant radio flares. In particular, transient gamma-ray emission was found when the system is moving into or out of the quenched state, a spectral state – characterized by a very low (or undetectable) flux at radio and hard X-ray frequencies – that is known to occur a few days before major radio flares.
The quenching activity of Cyg X-3 turns out to be a key condition for the observed activity above 100 MeV. According to theoretical models, a simple leptonic scenario – based on inverse Compton (IC) scattering between electrons/positrons accelerated in the jet and seed photons from the WR companion – can account for the flaring gamma-ray fluxes and the 4.8 h modulation detected by Fermi-LAT during the transient activity (Dubus et al., 2010). A simple phenomenological picture, based on dominant leptonic processes in the jet, can account for the non-thermal emission pattern: around the quenching, the jet would consist of plasmoids, ejected with high Lorentz factor. This transient jet would be responsible for the HE flare (for IC processes), produced in the proximity of the binary system ( cm), and it would subsequently produce the major radio flares (synchrotron processes), by moving out from the central engine (distances cm). MAGIC repeatedly observed Cyg X-3, both during hard and soft spectral states, but never detected any significant VHE activity from the microquasar (Aleksić et al., 2010).
@@
We carried out simulations by assuming two different theoretical models based on IC processes in the jet (Piano et al., 2012; Zdziarski et al., 2018), in order to test the possibility of a CTA detection of transient VHE gamma-rays from Cyg X-3.
@@
, by simulating observations with the Alpha Configuration of the Northern array of the CTA observatory (IRF: North_z20
), taking into account the energy dispersion. A multi-source input model with the main background TeV sources (see Section 4.1) and the CTA instrumental background (CTAIrfBackground
@@
In the first case, we adopted a simple power-law spectrum (see Eq. 1) inferred from the leptonic Model A from Piano et al. (2012), where prefactor ph , index and pivot energy TeV. The leptonic model is based on IC scatterings between accelerated electrons in the jet and soft seed photons from the accretion disk (X-rays) and from the companion star (UV). We simulated 5 hour and 50 hour observations, and we investigated the resulting simulated data, by performing a binned analysis. The resulting spectra are shown in Fig. 7, together with the X-ray “hypersoft” spectrum (Koljonen et al., 2010), the AGILE flaring spectrum (Piano et al., 2012) and the MAGIC flux ULs observed during the soft states (Aleksić et al., 2010). All the spectra (not simultaneous observations) are referred to the same spectral state of Cyg X-3 when the transient gamma-ray activity is detected at MeV-GeV energies (quenched state). We show the reference theoretical model and the input simulated power-law, together with the CTA simulated spectra. By assuming this input spectrum we found no detection with CTA-N with 5-h observation and a weak hint of signal (3) for 50 hours of observation time.
@@
In the second case, we assumed a different theoretical model, developed by Zdziarski et al. (2018) in order to fit the flaring spectrum from Cyg X-3 as detected by Fermi-LAT (cumulative spectrum of 49 1-day flares detected between August 2008 and August 2017). The theoretical model presented in their paper is similar to the one presented in Piano et al. (2012), but in Zdziarski et al. (2018), the electrons in the jet scatter blackbody soft photons from the companion star only. The orbital and geometrical parameters are similar. Also in this case, the model is focused on the HE emission from the microquasar (E 100 GeV). Thus, we assumed a simple power-law extension of the model up to TeV energies (assuming: prefactor ph , index and pivot energy TeV). Similarly, we simulated 5 h and 50 h observations. The results of these simulations are shown in Fig. 8. In this case, by assuming a harder and brighter input spectrum, we found clear detections with CTA-N: 10 with 5-h observation, 30 with-50 h observation.
@@
Thus, by assuming two simple power-law input spectra adapted from theoretical leptonic models – both created ad hoc in order to account for the flaring activity observed by AGILE and Fermi-LAT – a possible detection with CTA North is plausible even with a few hours observations. We cannot rule out to detect the 4.8 h orbital modulation, in the case of a prolonged TeV flare. A CTA detection of transient VHE gamma-ray activity would represent an unprecedented result for this elusive system, never observed at TeV energies. Nevertheless, a CTA non-detection would give new strong constraints on theoretical models about microquasars. The lack of a transient VHE signal from Cyg X-3, correlated with non-thermal flaring activity, could indicate that: (i) the TeV signal, eventually produced in the jet, is absorbed for pair production by the companion star’s UV photons; (ii) the acceleration efficiency in the jet is intrinsically low, the maximum energies of the jet particles are not sufficient to generate TeV photons.
@@ 4.1.3 Cyg X-1
Cyg X-1 is a HMXB, composed of a black hole ( ) and a O9.7Iab supergiant companion star ( , Miller-Jones et al., 2021). The system is located at a distance of kpc (Miller-Jones et al., 2021), and the orbital period is 5.6 days. The X-ray spectra can be accurately modeled by hybrid Componization models (Coppi, 1999). The soft state of Cyg X-1 is characterized by a strong disk blackbody component peaking at keV and a power-law tail extending up to 10 MeV, related to Componization processes in the corona. In the hard state, the accretion disk is truncated and the emission from the corona is dominant. In this state, the coronal plasma is composed by a hot quasi-thermal population of electrons ( keV) with a sharp cutoff at 200 keV. At sub-MeV energies, the microquasar exhibits a non-thermal power-law tail with a strong linear polarization (Laurent et al., 2011; Jourdain et al., 2012). This emission could be ascribed either to synchrotron processes in the jet, by assuming a very efficient particle acceleration and strong jet magnetic fields (Zdziarski et al., 2014), or to the corona itself (Romero et al., 2014). Recent studies investigate the physical origin of this power-law tail at sub-MeV energies, detected during both soft and hard spectral states (Cangemi et al., 2021). Above 100 MeV, deep observations with Fermi-LAT found evidences of persistent emission from Cyg X-1 only during hard X-ray spectral states (Zanin et al., 2016; Zdziarski et al., 2017). Transient HE emission was observed by AGILE (Bulgarelli et al., 2010; Sabatini et al., 2010, 2013) on 1-2 day timescales, in coincidence with both hard and soft X-ray spectral states. At TeV energies, a hint of detection (4) was observed by MAGIC on September 24, 2006 (Albert et al., 2007), during a hard X-ray flare of Cyg X-1.
For Cyg X-1, we investigated the possibility that CTA will detect both transient and persistent emission from the microquasar.
@@
In this case, we carried out a simulated short-term observation of Cyg X-1, during a possible VHE gamma-ray flare. We simulated a 30-minute observation with the same setup reported in Section 4.1.2: a multi-source simulation with photon energies between 100 GeV and 1 TeV. We assumed, as input spectrum for the simulation, the same power-law observed by MAGIC in September 24, 2006 (Albert et al. (2007); prefactor ph , index , pivot energy TeV). We obtained an overall detection of the source at a significance level of 38. The resulting spectrum is shown in Fig. 9, together with the observed flaring spectrum observed by MAGIC. Our results confirm that CTA will be able to detect a flare similar to the one reported by MAGIC in 2006 in a few minute observation, with unprecedented spectral accuracy.
@@
Cyg X-1 exhibits persistent HE emission during the hard state, as observed by Fermi-LAT (Zanin et al., 2016; Zdziarski et al., 2017). Thus, we investigated the possibility of a CTA detection of VHE persistent emission above 100 GeV. Again, we assumed the same setup as reported in Section 4.1.2. We analyzed three different scenarios. In the first one, we assumed as input spectral model for CTA, a simple extension of the power-law spectral shape reported in the Fermi-LAT 4FGL Catalog, without any cut off around 100 GeV (Abdollahi et al., 2020). In the second scenario, we assumed a spectral shape based on a purely leptonic theoretical model, in which gamma-ray emission is produced due to IC scatterings in the persistent jet during the hard state (Zdziarski et al., 2017). According to this model, a sharp cut off – due to the Klein-Nishina effects – is predicted at 100 GeV. In the third scenario, we assumed a spectral shape based on the lepto-hadronic theoretical model presented in Kantzas et al. (2021). In that paper, the authors modeled the GeV persistent spectrum as detected by Fermi-LAT during the hard state, by assuming that both electrons and protons are accelerated in the jet. A comprehensive model, based on a superposition of leptonic (IC scatterings) and hadronic processes (gamma rays from the decay of neutral mesons, produced in interactions) can properly fit the multi-wavelength spectrum up to the high-energy emission from Cyg X-1.
@@
For the second hypothesis (spectrum inferred from Zdziarski et al., 2017), we assumed a simple power-law (with prefactor ph , index and pivot energy TeV). A multi-source simulation with photon energies between 100 GeV and 1 TeV has been carried out. In this case, we did not detect any significant emission with CTA with a simulated observation of 50 h (significance 2). The resulting differential spectral ULs are shown in Fig. 11, together with the theoretical model from Zdziarski et al. (2017), the HE gamma-ray spectra as detected by Fermi-LAT (Zanin et al., 2016; Zdziarski et al., 2017), and the MAGIC ULs related to the hard state (Ahnen et al., 2017b).
For the third hypothesis (spectrum inferred from Kantzas et al., 2021), we used as input the theoretical model itself, by simulating photon energies between 100 GeV and 100 TeV. We carried out the usual multi-source binned analysis, and we found a clear detection with a significance of 36 for a 50h simulated observation. The resulting simulated spectrum is shown in Fig. 12, together with the theoretical model from Kantzas et al. (2021), the HE gamma-ray spectra as detected by Fermi-LAT (Zanin et al., 2016; Zdziarski et al., 2017), and the MAGIC flux ULs during the hard state (Ahnen et al., 2017b).
Thus, according to our simulations, CTA will be able to detect a possible persistent VHE gamma-ray emission from the jet of Cyg X-1, if the spectrum is not characterized by a sharp cut off around 100 GeV. According to purely leptonic models, a sharp cut off is expected below 1 TeV. On the contrary, hadronic processes could be responsible for a bright emission above 1 TeV, which could be detected by the CTA Observatory.

### v2

@@ 4.1.1 SS 433
Source | [TeV-1cm-2s | EB1 [TeV] | EB2 [TeV] | |||||
Central Source | ||||||||
East 1 | ||||||||
East 2 | ||||||||
East 3 | ||||||||
West 1 | ||||||||
West 2 | ||||||||
West 3 |
At TeV energies SS 433 was detected by both the High Altitude Water Cherenkov (HAWC) Observatory (1017 days of measurements, Abeysekara et al. (2018a)) and H.E.S.S. (200 hours of observations, H. E. S. S. Collaboration et al. (2024)). These observations demonstrate presence of two regions of gamma-ray emission of leptonic nature at the positions of the eastern and western jets. The reported H.E.S.S. fluxes at 1 TeV (( TeV-1 cm-2 s-1 and ( TeV-1 cm-2 s-1 for the eastern and western jets correspondingly) are inline with HAWC data. Quality of the H.E.S.S. data also allow to study the energy dependence of the source morphology, demonstrating that while the gamma-ray emission above 10 TeV appears only at the base of the jets, the lower-energy gamma rays have their peak surface brightness at locations further along each jet, reflecting an energy-dependent particle energy loss timescale.
@@
While we do not expect to detect variability in the VHE emission coming from the lobes, microquasars are known to have flaring emission on various timescales coming from its central source. To test the possibility of CTAO to detect a central source and its putative variability we simulate the local region of SS 433 with the diffuse background and the nearby MGRO 1906+06 source, where SS 433 consists of both the aforementioned lobes and a central point source.
Following the H.E.S.S. observations, we have modelled the eastern and western lobes as a combination of three hotspots with Gaussian profiles (with the parameters summarised in Table S4 of H. E. S. S. Collaboration et al. (2024)). Spectral models of the lobes were organised so that different hotspots appear in different energy bands. To represent the spectral model of each hotspot in agreement with H. E. S. S. Collaboration et al. (2024), we assumed that they follow a powerlaw distribution with a super-exponential cut-off at both high and low energies, to allow hotspots to arise at different energies:
The parameters chosen to represent the H.E.S.S. spectrum are given in Table 4. No low-energy cut-off was assumed for the e1 and w1 sources. The spectral model of the central source was obtained directly from the H.E.S.S. ULs, and was modelled using a simple power law model with a flux falling below the H.E.S.S. UL value. As it is seen in Figure 8, 20 hours of CTAO observations is enough to clearly measure the energy dependent source structure as well as to detect the central source at the assumed flux level. The dependence of the central source relative flux errors on the exposure time is shown in Figure 9.
To study the CTAO possibility to detect possible variability of about 15% with the precession and orbital periods at the level proposed by Rasul et al. (2019) we have simulated a 500 hours observation of the source uniformly distributed along the precessional period, assuming and 500 hours observation of the source uniformly distributed along the orbital period, assuming . The expected variability is shown in Figure 10.
To systematically study the CTAO sensitivity to detect different level of variability for various levels of the source flux we run batches of 5000 simulations for low exposure times of 30 mins and 60 mins, and batches of 1000 simulations for higher exposure times of 300, 600 and 6000 mins using the North and South site IRFs The flux was calculated after every simulation and the total data was compiled into histograms for each exposure time to determine the error range of the detections. As SS 433 could be viewed from both the north and south hemispheres the results of error measurements from both sites were compared to determine which can produce more sensitive detections. Figure 11 shows the comparison of this flux error ratio depending on the integrated source flux above 1 TeV and exposure for both Northern (left) and Southern arrays. In this figure one can see also simulations for the values exceeding the H.E.S.S. upper limit on the central source ( ph cm-2 s-1). This was done to test the effectiveness of CTAO on dim transient sources to determine what level of variability can be observed with short exposures.
@@
Based on the simulations run at different fluxes, sources with a photon flux ph cm-2 s-1 will require exposure times of more than 10 hours in order to detect flux variability at about 50% level. However, for sources with a flux ph cm-2 s-1, CTAO may be able to detect variability as low as observing from 5 to 10 hours. At a photon flux ph cm-2 s-1 the 1 sigma ratio gets as low as with exposure times of an hour long, meaning that even low variability may be detectable from relatively bright sources with short observations.
To be sure that our results are applicable to the sources located in crowded regions we compared our results assuming that the source with the flux of ph cm-2 s-1 was located either in the uncrowded region, like SS 433, or in the region with multiple nearby TeV sources, like LS 5039. It was found that the observed flux of the central source and its error agreed in these two cases within few percents and thus our results are valid for sources located in both crowded and uncrowded regions.
@@ 4.1.2 Cyg X-3
Cyg X-3 is a HMXB located at a distance of 9 kpc (Reid & Miller-Jones, 2023). The companion star is a Wolf-Rayet (WR) with a strong wind mainly composed of helium. The nature of the compact object is still unknown, although a black hole scenario is favored (Zdziarski et al., 2013; Antokhin et al., 2022). The orbital period is very short, hours, indicating that the compact object is very close to the WR star, totally enshrouded in its stellar wind (orbital distance cm). Recent observations with the Imaging X-ray Polarimetry Explorer (IXPE) show a high X-ray polarization degree from the system, during different spectral states, indicating the presence of collimated optically-thick outflows, which hide the central engine (Veledina et al., 2024a, b). The binary system is known to produce giant radio flares (flux > 10 Jy), produced by synchrotron processes from a relativistic jet oriented very close to the line of sight. Transient gamma-ray activity above 100 MeV was reported for the first time in 2009 by AGILE (Tavani et al., 2009) and Fermi-LAT (Fermi LAT Collaboration et al., 2009), and reported in several studies over the years, since its discovery (see Prokhorov & Moraghan, 2023 for a recent study on the transient activities observed by Fermi-LAT). The flaring activity (typical duration: 1-2 days) was observed in coincidence with a repetitive pattern of multi-frequency emission (Piano et al., 2012): the gamma-ray flares have been detected (i) during soft X-ray spectral states (around minima of the hard X-ray light curve), (ii) in the proximity of spectral transitions, and (iii) a few days before giant radio flares. In particular, transient gamma-ray emission was found when the system is moving into or out of the quenched state, a spectral state – characterized by a very low (or undetectable) flux at radio and hard X-ray frequencies – that is known to occur a few days before major radio flares.
The quenching activity of Cyg X-3 turns out to be a key condition for the observed activity above 100 MeV. According to theoretical models, a simple leptonic scenario – based on inverse Compton (IC) scattering between electrons/positrons accelerated in the jet and seed photons from the WR companion – can account for the flaring gamma-ray fluxes and the 4.8 h modulation detected by Fermi-LAT during the transient activity (Dubus et al., 2010; Prokhorov & Moraghan, 2023). A simple phenomenological picture, based on dominant leptonic processes in the jet, can account for the non-thermal emission pattern: around the quenching, the jet would consist of plasmoids, ejected with high Lorentz factor. This transient jet would be responsible for the HE flare (for IC processes), produced in the proximity of the binary system ( cm), and it would subsequently produce the major radio flares (synchrotron processes), by moving out from the central engine (distances cm). MAGIC repeatedly observed Cyg X-3, both during hard and soft spectral states, but never detected any significant VHE activity from the microquasar (Aleksić et al., 2010).
@@
We carried out simulations by assuming two different theoretical models based on IC processes in the jet (Piano et al., 2012; Zdziarski et al., 2018), in order to test the possibility of a CTAO detection of transient VHE gamma rays from Cyg X-3.
@@
, by simulating observations with the Alpha Configuration of the Northern array of the CTAO observatory (IRF: North_z20
), taking into account the energy dispersion. A multi-source input model with the main background TeV sources (see Section 4.1) and the CTAO instrumental background (CTAOIrfBackground
@@
In the first case, we adopted a simple power-law spectrum (see Eq. 1) inferred from the leptonic Model A from Piano et al. (2012), where prefactor ph , index and pivot energy TeV. The leptonic model is based on IC scatterings between accelerated electrons in the jet and soft seed photons from the accretion disk (X-rays) and from the companion star (UV). We simulated 5 hour and 50 hour observations, and we investigated the resulting simulated data, by performing a binned analysis. The resulting spectra are shown in Fig. 12, together with the X-ray “hypersoft” spectrum (Koljonen et al., 2010), the AGILE flaring spectrum (Piano et al., 2012) and the MAGIC flux ULs observed during the soft states (Aleksić et al., 2010). All the spectra (not simultaneous observations) are referred to the same spectral state of Cyg X-3 when the transient gamma-ray activity is detected at MeV-GeV energies (quenched state). We show the reference theoretical model and the input simulated power-law, together with the CTAO simulated spectra. By assuming this input spectrum we found no detection with CTAO-N with 5-h observation and a weak hint of signal (3) for 50 hours of observation time.
@@
In the second case, we assumed a different theoretical model, developed by Zdziarski et al. (2018) in order to fit the flaring spectrum from Cyg X-3 as detected by Fermi-LAT (cumulative spectrum of 49 1-day flares detected between August 2008 and August 2017). The theoretical model presented in their paper is similar to the one presented in Piano et al. (2012), but in Zdziarski et al. (2018), the electrons in the jet scatter blackbody soft photons from the companion star only. The orbital and geometrical parameters are similar. Also in this case, the model is focused on the HE emission from the microquasar (E 100 GeV). Thus, we assumed a simple power-law extension of the model up to TeV energies (assuming: prefactor ph , index and pivot energy TeV). Similarly, we simulated 5 h and 50 h observations. The results of these simulations are shown in Fig. 13. In this case, by assuming a harder and brighter input spectrum, we found clear detections with CTAO-N: 10 with 5-h observation, 30 with-50 h observation.
@@
Thus, by assuming two simple power-law input spectra adapted from theoretical leptonic models – both created ad hoc in order to account for the flaring activity observed by AGILE and Fermi-LAT – a possible detection with CTAO North is plausible even with a few hours observations. It is important to note that these extrapolation do not take into account absorption for pair production in the companion star’s photon field, which could be not negligible between 100 GeV and 1 TeV (Zdziarski et al., 2012). Nevertheless, we cannot rule out to detect the 4.8 h orbital modulation, in the case of a prolonged TeV flare. A CTAO detection of transient VHE gamma-ray activity would represent an unprecedented result for this elusive system, never observed at TeV energies. Nevertheless, a CTAO non-detection would give new strong constraints on theoretical models about microquasars. The lack of a transient VHE signal from Cyg X-3, correlated with non-thermal flaring activity, could indicate that: (i) the TeV signal, eventually produced in the jet, is absorbed for pair production by the companion star’s UV photons; (ii) the acceleration efficiency in the jet is intrinsically low, the maximum energies of the jet particles are not sufficient to generate TeV photons.
@@ 4.1.3 Cyg X-1
Cyg X-1 is a HMXB, composed of a black hole ( ) and a O9.7Iab supergiant companion star ( , Miller-Jones et al., 2021). The system is located at a distance of kpc (Miller-Jones et al., 2021), and the orbital period is 5.6 days. The X-ray spectra can be accurately modeled by hybrid Componization models (Coppi, 1999). The soft state of Cyg X-1 is characterized by a strong disk blackbody component peaking at keV and a power-law tail extending up to 10 MeV, related to Componization processes in the corona. In the hard state, the accretion disk is truncated and the emission from the corona is dominant. In this state, the coronal plasma is composed by a hot quasi-thermal population of electrons ( keV) with a sharp cutoff at 200 keV. At sub-MeV energies, the microquasar exhibits a non-thermal power-law tail with a strong linear polarization (Laurent et al., 2011; Jourdain et al., 2012). This emission could be ascribed either to synchrotron processes in the jet, by assuming a very efficient particle acceleration and strong jet magnetic fields (Zdziarski et al., 2014), or to the corona itself (Romero et al., 2014). Recent studies investigate the physical origin of this power-law tail at sub-MeV energies, detected during both soft and hard spectral states (Cangemi et al., 2021). Above 100 MeV, deep observations with Fermi-LAT found evidences of persistent emission from Cyg X-1 only during hard X-ray spectral states (Zanin et al., 2016; Zdziarski et al., 2017). Transient HE emission was observed by AGILE (Bulgarelli et al., 2010; Sabatini et al., 2010, 2013) on 1-2 day timescales, in coincidence with both hard and soft X-ray spectral states. At TeV energies, a hint of detection (4) was observed by MAGIC on September 24, 2006 (Albert et al., 2007), during a hard X-ray flare of Cyg X-1. A 4 persistent TeV signal was recently reported by LHAASO (LHAASO Collaboration, 2024).
For Cyg X-1, we investigated the possibility that CTAO will detect both transient and persistent emission from the microquasar.
@@
In this case, we carried out a simulated short-term observation of Cyg X-1, during a possible VHE gamma-ray flare. We simulated a 30-minute observation with the same setup reported in Section 4.1.2: a multi-source simulation with photon energies between 100 GeV and 1 TeV. We assumed, as input spectrum for the simulation, the same power-law observed by MAGIC in September 24, 2006 (Albert et al. (2007); prefactor ph , index , pivot energy TeV). We obtained an overall detection of the source at a significance level of 38. The resulting spectrum is shown in Fig. 14, together with the observed flaring spectrum observed by MAGIC. Our results confirm that CTAO will be able to detect a flare similar to the one reported by MAGIC in 2006 in a few minute observation, with unprecedented spectral accuracy. A fainter TeV flare - weaker than the one reported by MAGIC - would require a longer CTAO observation (a few hours) to be significantly detected. This possibility will be properly assessed in a potential ToO observation, on the basis of the triggering flare flux in other wavelengths.
@@
Cyg X-1 exhibits persistent HE emission during the hard state, as observed by Fermi-LAT (Zanin et al., 2016; Zdziarski et al., 2017). Thus, we investigated the possibility of a CTAO detection of VHE persistent emission above 100 GeV. Again, we assumed the same setup as reported in Section 4.1.2. We analyzed three different scenarios. In the first one, we assumed as input spectral model for CTAO, a simple extension of the power-law spectral shape reported in the Fermi-LAT 4FGL Catalog, without any cut off around 100 GeV (Abdollahi et al., 2020). In the second scenario, we assumed a spectral shape based on a purely leptonic theoretical model, in which gamma-ray emission is produced due to IC scatterings in the persistent jet during the hard state (Zdziarski et al., 2017). According to this model, a sharp cut off – due to the Klein-Nishina effects – is predicted at 100 GeV. In the third scenario, we assumed a spectral shape based on the lepto-hadronic theoretical model presented in Kantzas et al. (2021). In that paper, the authors modeled the GeV persistent spectrum as detected by Fermi-LAT during the hard state, by assuming that both electrons and protons are accelerated in the jet. A comprehensive model, based on a superposition of leptonic (IC scatterings) and hadronic processes (gamma rays from the decay of neutral mesons, produced in interactions) can properly fit the multi-wavelength spectrum up to the high-energy emission from Cyg X-1.
@@
For the second hypothesis (spectrum inferred from Zdziarski et al., 2017), we assumed a simple power-law (with prefactor ph , index and pivot energy TeV). A multi-source simulation with photon energies between 100 GeV and 1 TeV has been carried out. In this case, we did not detect any significant emission with CTAO with a simulated observation of 50 h (significance 2). The resulting differential spectral ULs are shown in Fig. 16, together with the theoretical model from Zdziarski et al. (2017), the HE gamma-ray spectra as detected by Fermi-LAT (Zanin et al., 2016; Zdziarski et al., 2017), and the MAGIC ULs related to the hard state (Ahnen et al., 2017b).
For the third hypothesis (spectrum inferred from Kantzas et al., 2021), we used as input the theoretical model itself, by simulating photon energies between 100 GeV and 100 TeV. We carried out the usual multi-source binned analysis, and we found a clear detection with a significance of 36 for a 50h simulated observation. The resulting simulated spectrum is shown in Fig. 17, together with the theoretical model from Kantzas et al. (2021), the HE gamma-ray spectra as detected by Fermi-LAT (Zanin et al., 2016; Zdziarski et al., 2017), the MAGIC flux ULs during the hard state (Ahnen et al., 2017b) and the HAWC flux ULs between 0.1 and 100.0 TeV (Albert et al., 2021). In particular, we note that above 10 TeV the HAWC ULs for a prolonged stacked observation (1523 days) are below our theoretical model. This spectral behavior could weaken the chance of a CTAO sharp detection at the highest energies.
Thus, according to our simulations, CTAO will be able to detect a possible persistent VHE gamma-ray emission from the jet of Cyg X-1, if the spectrum is not characterized by a sharp cut off around 100 GeV. According to purely leptonic models, a sharp cut off is expected below 1 TeV. On the contrary, hadronic processes could be responsible for a bright emission above 1 TeV, which could be detected by the CTAO Observatory.